# Aligning Attention Distribution to Information Flow for Hallucination Mitigation in Large Vision-Language Models

### v1

@@ 4.2 Mitigating Aggravated Language priors
We optimize the model’s attention allocation to align with the flow pattern of visual information, which ultimately results in increased attention to semantic representations. However, while semantic representations incorporate visual information, they also inherently contain their own contextual features. As a result, increasing the model’s reliance on semantic representations will exacerbate the issue of language priors Leng et al. (2024); Wu et al. (2022).
To address this issue, we use the CICD method Zhao et al. (2025) to eliminate language priors while preserving visual information. CICD uses the cross-images consistency of language priors to identify detrimental priors and mitigate them by contrastive decoding:
@@
where $\hat{\mathrm{logit}}_{\theta}$ is the logit distribution after attention allocation alignment and $\mathrm{logit}_{\theta}$ is the regular logit distribution with a distinct image $v^{\prime}$. $\alpha$ is defined as follows:
@@ 5 Experiments
VCD 52.91 55.7 16.8 53.20 59.6 18.9
ICD 52.82 53.9 16.6 53.24 55.7 16.7
IBD 52.48 54.1 15.8 54.14 56.4 15.5
VAR 52.98 55.1 17.1 52.63 55.9 16.9
VAF 52.36 55.6 18.1 52.63 55.9 16.9
DeFG 52.72 55.7 16.6 53.06 59.1 17.4
CICD 55.80 45.6 13.1 54.20 45.7 13.2
@@
VCD 9.8 51.2 43.8 4.4 9.9 54.0 44.6 4.2
ICD 8.8 51.2 38.7 4.1 9.8 53.9 46.7 5.1
IBD 9.8 50.5 42.2 4.4 9.0 56.1 45.1 4.6
VAR 11.7 51.2 48.5 4.8 - - - -
VAF 11.3 50.2 48.6 4.3 11.5 51.8 50.1 5.1
DeFG 9.1 50.7 39.9 4.1 9.7 54.1 44.5 5.2
CICD 6.6 52.7 34.8 2.2 7.1 53.6 35.0 2.3
SEVIBalanced 5.6 48.6 27.6 1.7 6.0 51 28.5 1.6
@@
5.1 Experiment Setting
Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec. A.
@@ 5.1 Experiment Setting
Our method guides the model to focus on core information, thereby reducing the occurrence of hallucinations. In practice, we observe a trade-off between hallucination suppression and the richness of generated details, influenced by the model’s degree of attentional focus. To address this, we introduce hyperparameters to control the model’s focus on core information. Based on empirical observations, we design two sets of hyperparameter configurations: (1) Focused mode that aggressively minimizes hallucinations, and (2) Balanced mode that strikes a compromise between detail retention and hallucination reduction. Specifically, the Focused mode optimizes attention distribution starting from the 5th layer with $\omega=4$, while the Balanced mode begins from the 9th layer with $\omega=0.5$.
In addition, we set the parameter $\kappa$ in Eq. 2 to 0.2. Further details are discussed in Appendix. A.3. We employ sampling decoding for the next-token prediction with default settings. All experiments are performed on a single NVIDIA A800 40G GPU.
@@
We evaluate the performance of our method across three widely adopted multimodal hallucination benchmarks on the image captioning task. These include CHAIR Rohrbach et al. (2018), DetailCaps Ye et al. (2025), and AMBER Wang et al. (2024b).
@@
To examine the generalizability of our approach, we apply it to five LVLMs drawn from three representative model families: InstructBLIP Dai et al. (2023); two models from the LLaVA family (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)); and two from the Qwen series (Qwen2-VL-Instruct Wang et al. (2024c) and Qwen2.5-VL-Instruct Bai et al. (2025)). All models are tested at the 7B parameter scale unless explicitly noted otherwise.
@@
We conduct a comparison between our method and several SOTA de-hallucination techniques: VCD Leng et al. (2024), ICD Wang et al. (2024d), IBD Zhu et al. (2024), VAR Kang et al. (2025), VAF Yin et al. (2025), DeFG Zhang et al. (2025), and CICD Zhao et al. (2025).
5.2 Main Results
Results on CHAIR
CHAIR is a benchmark designed to detect object hallucinations in image captions, relying on human annotations to provide reliable ground truth. Following common practice, we conduct experiments with maximum sequence lengths set to 64 and 512, respectively. The results are shown in Fig. 1. Our method effectively reduces hallucinations by guiding the model’s attention toward key information, achieving particularly low hallucination rates under the focused mode. However, we observe that when the model concentrates on critical information, it tends to become more conservative, as reflected by a slight drop in recall. In contrast, the balanced mode allows the model to capture more details while still maintaining a low hallucination rate, resulting in the best overall performance. In addition, our method yields the most pronounced improvement on the 13B model relative to the 7B model, indicating its effectiveness in harnessing the latent capacity of larger models for visual understanding.
Method LLaVA-next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
@@ 5.3 Effectiveness on more LVLMs
Setting LLaVA-1.5 InstructBLIP
[…]
Regular 54.6 16.4 62.6 19.5
w/o CICD 38.6 13.9 44.4 20.3
w/o Core Heads 39.0 9.7 45 13.3
w/o Global Heads 22.2 7.0 23.2 11.6
SEVIFocused 17.8 5.5 18.8 8.4
Table 5: Ablation study on CHAIR with Focused mode.
@@ 5.4 Ablation Study
Through a two-stage optimization paradigm, we reallocate the model’s attention toward key information, encouraging it to focus on the core content of the image and thereby reducing hallucinations. In addition, we employ the CICD method to address the issue of language priors that aggravate when the model over-focuses on semantic representations. We conduct ablation studies on these three components, with the results presented in Tab. 5. The core semantic heads play a pivotal role in guiding the model’s focus, and removing them in the ablation study leads to a noticeable increase in hallucination rate. The global semantic heads help modulate the cross-modal attention distribution, aligning it with the flow of visual information, which also contributes significantly to hallucination reduction. The combination of both components has a synergistic effect, and incorporating the CICD method to mitigate language priors, amplified by overattending to semantic representations, further enhances the performance of the model. The experimental results confirm the effectiveness and soundness of our method’s design, highlighting the value of each component.
@@ 6 Conclusion
7 Limitation
Our method guides the model to focus on the most critical information, thereby reducing the occurrence of hallucinations. During our experiments, we observed a trade-off mechanism in the model’s focusing process: excessive focus leads to more conservative image captioning. In an effort to avoid hallucinated content, the model may overlook some details, which is reflected in a slight decrease in recall. Although our method demonstrates superior overall performance, showing a significant advantage on comprehensive metrics such as CAPTURE, it still faces a trade-off between reducing hallucinations and generating more details. To address this issue, we introduce hyperparameters to control the model’s level of conservativeness, allowing users to manually adjust the behavior based on specific application scenarios. Furthermore, as a training-free approach, our method offers greater usability but is inherently limited by the performance ceiling of the model itself.
Appendix A Detailed Experiments
A.1 Benchmarks
@@
CHAIRs=|Hallucinated Objects||All $\text{CHAIRs}=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (7)
CHAIRi=|Hallucinated Sentences||All $\text{CHAIRi}=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
@@
Wang et al. (2024b) contains 1,004 carefully curated high-quality images, each with manually annotated objects. AMBER contains multiple evaluation metrics: CHAIR, Cover, Hal, and Cog. Given a list of annotated objects $A_{obj}={obj_{1}^{A},obj_{2}^{A},\cdots,obj_{n}^{A}}$ and a set of generated objects $R^{\prime}_{obj}$, each metric is defined as follows:
@@ A.2 Baselines
  • ICD Wang et al. (2024d) constructs negative contexts by designing adversarial instructions and applies contrastive decoding to mitigate their influence. Like VCD, it faces challenges such as visual information loss and performance bias.
  • IBD Zhu et al. (2024) strengthens the model’s focus on visual information by using the original context as a negative reference, and further refines contrastive decoding via inter-layer and inter-context consistency mechanisms.
@@
  • DeFG Zhang et al. (2025) generates negative contexts via cross-modal back-translation and strengthens consistency signals throughout the contrastive decoding process.
  • CICD Zhao et al. (2025) leverages the consistency of language priors across images by using different images to construct negative contexts. It detects harmful priors via consistency analysis and removes them through contrastive decoding, while retaining beneficial priors essential for accurate understanding.
@@ A.3 Hyperparameters
Appendix B Case Study

### v3

@@ 4.2 Mitigating Aggravated Language Priors
We optimize the model’s attention distribution to align with the flow pattern of visual information, which ultimately results in increased attention to semantic representations. However, while semantic representations incorporate visual information, they also inherently contain their own contextual features. As a result, increasing the model’s reliance on semantic representations will exacerbate the issue of language priors Wu et al. (2022); Ren et al. (2023); Chen et al. (2025).
To address this issue, we employ CICD Zhao et al. (2025) to eliminate language priors while preserving visual information. CICD uses the cross-image consistency of language priors to identify detrimental priors and mitigate them by contrastive decoding:
@@
where $\hat{\mathrm{logit}}_{\theta}$ is the logit distribution after attention distribution alignment and $\mathrm{logit}_{\theta}$ is the regular logit distribution with a distinct image $v^{\prime}$. $\alpha$ is defined as follows:
@@ 5 Experiments
VCD Leng et al. (2024) 52.91 55.7 16.8 53.20 59.6 18.9
ICD Wang et al. (2024c) 52.82 53.9 16.6 53.24 55.7 16.7
PAI Liu et al. (2024e) 53.49 39.4 11.8 53.27 62.6 18.3
IBD Zhu et al. (2025a) 52.48 54.1 15.8 54.14 56.4 15.5
VAR Kang et al. (2025) 52.98 55.1 17.1 - - -
AD-HH Yang et al. (2025) 52.52 44.4 11.5 - - -
VAF Yin et al. (2025) 52.36 55.6 18.1 52.63 55.9 16.9
DeGF Zhang et al. (2025) 52.72 55.7 16.6 53.06 59.1 17.4
CICD Zhao et al. (2025) 55.80 45.6 13.1 54.20 45.7 13.2
@@
VCD Leng et al. (2024) 9.8 51.2 43.8 4.4 9.9 54.0 44.6 4.2
ICD Wang et al. (2024c) 8.8 51.2 38.7 4.1 9.8 53.9 46.7 5.1
PAI Liu et al. (2024e) 7.7 49.3 36.9 3.3 11.7 52.8 55.1 5.4
IBD Zhu et al. (2025a) 9.8 50.5 42.2 4.4 9.0 56.1 45.1 4.6
PAI Liu et al. (2024e) 7.7 49.3 36.9 3.3 11.7 52.8 55.1 5.4
VAR Kang et al. (2025) 11.7 51.2 48.5 4.8 - - - -
AD-HH Yang et al. (2025) 9.0 48.0 40.8 3.0 - - - -
VAF Yin et al. (2025) 11.3 50.2 48.6 4.3 11.5 51.8 50.1 5.1
DeGF Zhang et al. (2025) 9.1 50.7 39.9 4.1 9.7 54.1 44.5 5.2
CICD Zhao et al. (2025) 6.6 52.7 34.8 2.2 7.1 53.6 35.0 2.3
SEVIBalanced 5.6 48.6 27.6 1.7 6.0 51.0 28.5 1.6
@@
5.1 Experimental Setup
Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Appendix B.
@@ 5.1 Experimental Setup
Our method guides the model to focus on core information, thereby reducing the occurrence of hallucinations. In practice, we observe a trade-off between hallucination suppression and the richness of generated details, influenced by the model’s degree of attentional focus. To address this, we introduce hyperparameters to control the model’s focus on the core information. Based on empirical observations, we design two sets of hyperparameter configurations: (1) Focused mode that aggressively minimizes hallucinations, and (2) Balanced mode that strikes a compromise between detail retention and hallucination reduction. Specifically, Focused mode optimizes attention distribution starting from the 5th layer with $\omega=4$, while Balanced mode begins from the 9th layer with $\omega=0.5$.
In addition, we set the parameter $\kappa$ in Eq. 2 to 0.2. Further details are discussed in Appendix B.3. We employ sampling decoding for the next-token prediction with default settings. All experiments are performed on a single NVIDIA A800 40G GPU.
@@
We evaluate the performance of our method across three widely adopted multimodal hallucination benchmarks on the image captioning task. These include CHAIR Rohrbach et al. (2018), DetailCaps Ye et al. (2025), and AMBER Wang et al. (2023a).
@@
To examine the generalizability of our approach, we apply it to five LVLMs drawn from three representative model families: InstructBLIP Dai et al. (2023); two models from the LLaVA family (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)); and two from the Qwen series (Qwen2-VL-Instruct Wang et al. (2024b) and Qwen2.5-VL-Instruct Bai et al. (2025)). All models are tested at the 7B parameter scale unless explicitly noted otherwise.
@@
We conduct a comparison between our method and several SOTA de-hallucination techniques: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2025a), PAI Liu et al. (2024e), VAR Kang et al. (2025), VAF Yin et al. (2025), AD-HH Yang et al. (2025), DeGF Zhang et al. (2025), and CICD Zhao et al. (2025).
Method LLaVA-Next Qwen2-VL Qwen2.5-VL
@@
5.2 Main Results
Results on CHAIR
CHAIR is a benchmark designed to detect object hallucinations in image captions, relying on human annotations to provide reliable ground truth. Following common practice, we conduct experiments with maximum sequence lengths set to 64 and 512, respectively. The results are shown in Fig. 1. Our method effectively reduces hallucinations by guiding the model’s attention toward core information, achieving particularly low hallucination rates under the focused mode. However, we observe that when the model concentrates on critical information, it tends to become more conservative, as reflected by a slight drop in recall. In contrast, the balanced mode allows the model to capture more details while still maintaining a low hallucination rate, resulting in the best overall performance. In addition, our method yields the most pronounced improvement on the 13B model relative to the 7B model, indicating its effectiveness in harnessing the latent capacity of larger models for visual understanding.
Setting Cs $\downarrow$ Ci $\downarrow$ R $\uparrow$
PAI with different layers
None (only CD) 25.4 7.6 61.9
[[2-15]] (Lower) 20.8 6.8 56.7
[[16-31]] (Higher) 24.6 8.0 59.3
[[2-31]] (PAI) 20.0 6.0 57.6
SEVI combined with PAI
Regular 24.4 8.9 56.6
SEVIFocused 14.8 4.7 54.1
SEVIFocused w/ PAI 14.2 5.3 52.2
SEVIBalanced 18.8 5.5 59.5
SEVIBalanced w/ PAI 15.0 5.3 55.4
Table 5: PAI increase visual attention from the 2nd layer to the final layer ([[2-31]]). None (only CD) denotes the use of contrastive decoding alone, without increasing visual attention.
Setting LLaVA-1.5 InstructBLIP
$downarrow Cs\downarrow Ci\downarrow$
Regular 54.6 16.4 62.6 19.5
w/o CICD 38.6 13.9 44.4 20.3
w/o Core Heads 39.0 9.7 45.0 13.3
w/o Global Heads 22.2 7.0 23.2 11.6
w/o Two-Stage Opt. 18.4 9.6 35.2 10.7
SEVIFocused 17.8 5.5 18.8 8.4
Table 6: Ablation study on CHAIR with Focused mode. w/o Two-Stage Opt. represents applying core semantic heads to non-semantic heads. We find that this approach disrupts the model’s internal representations and leads to response collapse. To improve the smoothness of attention alignment, we set $\omega=1$ in this setting.
@@ 5.2 Main Results
5.3 Visual or Semantic Representations, Which Should LVLMs Focus On?
Some works Liu et al. (2024e); Zhu et al. (2025a) emphasize the importance of visual attention, whereas we encourage the model to focus more on semantic representations. To investigate whether the model should prioritize visual or semantic representations, we take PAI Liu et al. (2024e) as a contrast, which directly increases the model’s attention to visual representations.
As we analyze in Fig. 1 and Fig. 2, visual information resides in visual representations during the early layers of LVLMs and is gradually integrated into semantic representations in later layers. Therefore, we hypothesize that LVLMs should focus on visual representations in the lower layers and shift to semantic representations in the higher layers. We perform ablation studies on both lower and higher layers within the PAI to better understand the layer-wise impact. The results at the top of Tab. 5 are consistent with our hypothesis, indicating that it is more beneficial for the model to focus on visual representations in the lower layers. Furthermore, we explore combining PAI with our method by applying PAI to the lower layers (layers [2–15]) to enhance visual attention, while adopting our approach in the higher layers (layers [16–31]). The results at the bottom of Tab. 5 further demonstrate our hypothesis.
@@ 5.5 Ablation Study
Through a two-stage optimization paradigm, we reallocate the model’s attention toward the semantic representations, encouraging it to focus on the core visual information and thereby reducing hallucinations. In addition, we employ CICD to address the issue of language priors that aggravate when the model over-focuses on semantic representations. We conduct ablation studies on these three components, with the results presented in Tab. 6. The core semantic heads play a pivotal role in guiding the model’s focus, and removing them in the ablation study leads to a noticeable increase in hallucination rate. The global semantic heads help modulate the cross-modal attention distribution, aligning it with the flow of visual information, which also contributes significantly to hallucination reduction. The combination of both components has a synergistic effect, and incorporating the CICD method to mitigate language priors, amplified by overattending to semantic representations, further enhances the performance of the model. Additionally, the two-stage optimization paradigm also has a positive effect, making the optimization process more stable and smooth. The experimental results confirm the effectiveness and soundness of our method’s design, highlighting the value of each component.
@@ 6 Conclusion
Limitations
Our method guides the model to focus on the most critical information, thereby reducing the occurrence of hallucinations. During our experiments, we observed a trade-off mechanism in the model’s focusing process: excessive focus leads to a more conservative image captioning. In an effort to avoid hallucinated content, the model may overlook some details, which is reflected in a slight decrease in recall. Although our method demonstrates superior overall performance—showing a significant advantage on comprehensive metrics such as CAPTURE—it still faces a trade-off between reducing hallucinations and generating more details. To address this issue, we introduce hyperparameters to control the model’s level of conservativeness, allowing users to manually adjust the behavior based on specific application scenarios. Furthermore, as a training-free approach, our method offers greater usability but is inherently limited by the performance ceiling of the model itself.
Acknowledgments
We thank all reviewers for their detailed reviews and valuable comments. This work is supported by Zhongguancun Academy Project No.20240103.
Appendix A Two-Stage Optimization
We present the algorithm of our method in Algorithm 1.
Appendix B Detailed Experiments
B.1 Benchmarks
@@
$\text{CHAIRs}=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (8)
$\text{CHAIRi}=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
@@
Wang et al. (2023a) contains 1,004 carefully curated high-quality images, each with manually annotated objects. AMBER contains multiple evaluation metrics: CHAIR, Cover, Hal, and Cog. Given a list of annotated objects $A_{obj}={obj_{1}^{A},obj_{2}^{A},\cdots,obj_{n}^{A}}$ and a set of generated objects $R^{\prime}_{obj}$, each metric is defined as follows:
@@ B.2 Baselines
  • ICD Wang et al. (2024c) constructs negative contexts by designing adversarial instructions and applies contrastive decoding to mitigate their influence. Like VCD, it faces challenges such as visual information loss and performance bias.
  • IBD Zhu et al. (2025a) strengthens the model’s focus on visual information by using the original context as a negative reference, and further refines contrastive decoding via inter-layer and inter-context consistency mechanisms.
@@
  • DeGF Zhang et al. (2025) generates negative contexts via cross-modal back-translation and strengthens consistency signals throughout the contrastive decoding process.
  • CICD Zhao et al. (2025) leverages the consistency of language priors across images by using different images to construct negative contexts. It detects detrimental priors via consistency analysis and removes them through contrastive decoding, while retaining beneficial priors essential for accurate understanding.
Algorithm 1 Attention Distribution Alignment within a Layer
0: attention weights $\mathbf{W}\in\mathbb{R}^{H\times L\times L}$, the end indexes of visual tokens $e$, hyperparameters $\kappa$ and $\omega$
0: aligned attention weights $\hat{\mathbf{W}}$
$\mathbf{W}_{\text{cur}}\leftarrow\mathbf{W}[:,-1,:]$  # Attention weights of the last token $\mathbf{W}_{\text{sf}}\leftarrow\operatorname{Softmax}(\mathbf{W}_{\text{cur}})$  # Apply softmax normalization $\mathbf{W}_{S}\leftarrow\mathbf{W}_{\text{sf}}[:,e+1:]$  # Attention to semantic representations $\mathbf{W}_{O}\leftarrow\mathbf{W}_{\text{sf}}[:,:e+1]$  # Attention to other representations
 # Head categorization: $H_{S}\leftarrow(\sum\mathbf{W}_{S}>\sum\mathbf{W}_{O})$  # Semantic heads $H_{O}\leftarrow\neg H_{S}$  # Other heads $H_{c}\leftarrow(\max\mathbf{W}_{S}>\kappa\cdot\sum\mathbf{W}_{S})$  # Heads focusing on core information $H_{g}\leftarrow\neg H_{c}$  # Heads fail to focus on core information $H_{S_{g}}\leftarrow H_{S}\cap H_{g}$  # Global semantic heads $H_{S_{c}}\leftarrow H_{S}\cap H_{c}$  # Core semantic heads
 # Two-Stage Optimization:
if $|H_{O}|>0$ and $|H_{S_{g}}|>0$ then
  # Guide the other heads using the global semantic heads $m_{1}\leftarrow\frac{1}{|H_{S_{g}}|}\sum\mathbf{W}_{\text{cur}}[H_{S_{g}},:]$  # Average pooling $\mathbf{W}_{\text{cur}}[H_{O},:]\leftarrow(\mathbf{W}_{\text{cur}}[H_{O},:]+\omega\cdot m_{1})/(1+\omega)$
end if
if $|H_{S_{c}}|>0$ and $|H_{S_{g}}|>0$ then
  # Guide the global semantic heads using the core semantic heads $m_{2}\leftarrow\max\mathbf{W}_{\text{cur}}[H_{S_{c}},:]$  # Max pooling $\mathbf{W}_{\text{cur}}[H_{S_{g}},:]\leftarrow(\mathbf{W}_{\text{cur}}[H_{S_{g}},:]+\omega\cdot m_{2})/(1+\omega)$
end if
$\mathbf{W}[:,-1,:]\leftarrow\mathbf{W}_{\text{cur}}$  # Update attention weights
Return: $\mathbf{W}$
@@ B.3 Hyperparameters
Appendix C Case Studies